# Explaining higher-order correlations between elliptic and triangular flow

### v1

@@ Abstract
The ALICE Collaboration has analyzed a number of cumulants mixing elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), involving up to $8$ particles, in Pb+Pb collisions at the LHC. We unravel an unexpected simplicity in these complex mathematical quantities for collisions at fixed impact parameter. We show that as one increases the order in $v_{2}$, for a given order in $v_{3}$, the changes in the cumulants are solely determined by the mean elliptic flow in the reaction plane, which originates from the almond-shaped geometry of the overlap area between the colliding nuclei. We derive simple analytic relations between cumulants of different orders on this basis. Some of these relations are in reasonable agreement with existing data. We postulate that agreement will be much improved if the analysis is repeated with a finer centrality binning and a larger pseudorapidity acceptance.
@@ I Introduction
In this paper, we unravel the information contained in cumulants of azimuthal correlations involving both elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), which are called “mixed harmonic cumulants” (MHC) ALICE:2021adw . $v_{2}$ and $v_{3}$ are the largest Fourier harmonics of anisotropic flow ALICE:2011ab . Relativistic hydrodynamics predicts Niemi:2012aj that they are approximately proportional to the initial anisotropies $\varepsilon_{2}$ and $\varepsilon_{3}$, which are Fourier coefficients of the initial density profile. Therefore, cumulants involving $v_{2}$ and $v_{3}$ constrain models of the initial state more directly than higher harmonics, which are largely driven by nonlinear response terms Borghini:2005kd ; Gardim:2011xv . The lowest-order MHC is a 4-particle cumulant introduced by Bilandzic et al. Bilandzic:2013kga , which has been measured in Pb+Pb collisions at the LHC by ALICE ALICE:2016kpq and ATLAS ATLAS:2019peb . It represents the linear correlation between $v_{2}^{2}$ and $v_{3}^{2}$, which is negative except in ultra-central collisions. We have recently shown Alqahtani:2025wan that it is driven by the correlation between the elliptic flow in the reaction plane, $v_{2,x}$, and $v_{3}^{2}$. Our goal is to extend this study to higher-order cumulants, of 6 and 8 particles, which have subsequently been measured ALICE:2021adw .
@@
In Sec. II, we define the cumulants of $v_{2}$ and $v_{3}$ measured by ALICE using the formalism of generating functions and we briefly discuss the data. In Sec. III, we define similar cumulants but in a different coordinate frame, the “intrinsic” frame where the direction of impact parameter is fixed Roubertie:2025qps ; Alqahtani:2025wan . In this frame, the magnitude of cumulants decreases as a function of the order, in a way which follows specific scaling rules, provided that all events have the same impact parameter. In Sec. IV, we express the experimental cumulants as a function of the cumulants in the intrinsic frame. We explain the orders of magnitude of cumulants measured by ALICE. In Sec. V, we derive relations between cumulants of different orders, and compare them with experimental data.
@@ II Cumulants of flow fluctuations in the laboratory frame
The quantities that can be measured experimentally are moments of the joint distribution of $V_{2}$ and $V_{3}$ Bhalerao:2011yg ; Bhalerao:2014xra . The cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ analyzed by ALICE ALICE:2021adw , where $m$ and $q$ are positive integers, are combinations of these moments Bilandzic:2010jr ; Bilandzic:2013kga . [^1] The ALICE collaboration has subsequently analyzed a different set of “cumulants” which are different combination of moments ALICE:2023lwx . They are not derived from the same generating function as the cumulants studied in this paper, and our analysis does not apply to them. Their expressions are derived by expanding the generating function of cumulants Taghavi:2020gcy , which depends on complex variables Mehrabpour:2020wlu $\lambda$ and $\mu$ and their complex conjugates $\lambda^{*}$ and $\mu^{*}$:
@@
It was measured by ALICE in 2016 in Ref. ALICE:2016kpq , where it was named $SC(3,2)$. Higher-order cumulants with $(m,q)=(2,1),(3,1),(1,2),(2,2),(1,3)$ were subsequently measured in Ref. ALICE:2021adw , where their expressions in terms of moments are provided. Deriving these expressions is straightforward using Eq. (2). We do not repeat them because they are lengthy and will not be needed in this work. [^3] Note that the six MHC cumulants analyzed by ALICE only involve the magnitudes, not the phases of $V_{2}$ and $V_{3}$. The phases only enter at the next order $(m,q)=(3,2)$, a cumulant of order 10 whose expression involves moments of the type $\langle V_{2}^{3}(V_{3}^{*})^{2}\rangle$, and which has not yet been analyzed.
Figure 1: Comparison between ALICE ALICE:2021adw and ATLAS ATLAS:2019peb data on $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb+Pb collisions at 5.02 TeV per nucleon pair, as a function of the collision centrality. ATLAS denotes this quantity by $nsc_{2,3}\{4\}$.
ALICE normalizes the cumulants as follows:
@@
This normalization suppresses the sensitivity to kinematic cuts, and also provides an intuitive, dimensionless measure of the strength of the correlation. A value of order unity would imply a strong correlation between $v_{2}$ and $v_{3}$. But the measured values of $nMHC(v_{2}^{2m},v_{3}^{2q})$ are, in absolute magnitude, smaller than unity for all centralities. [^4] By contrast, $nMHC(v_{2}^{2},v_{4}^{2})$ reaches $1$ in peripheral collisions ATLAS:2019peb ; ALICE:2021adw , which is a natural consequence of the non-linear coupling between $v_{2}$ and $v_{4}$ Teaney:2012ke . They increase somewhat with $m$, and decrease strongly with $q$, which will be explained in Sec. IV.
The normalization (5) also facilitates the comparison with other experiments and theory calculations. A comparison between ALICE and ATLAS ATLAS:2019peb data for the lowest-order cumulant $nMHC(v_{2}^{2},v_{3}^{2})$ is displayed in Fig. 1. There are sizable differences between the two experiments, whose origin is unknown. One possible explanation is the wider centrality bins used by ALICE. However, one would typically expect wider bins to increase the value of $nMHC(v_{2}^{2},v_{3}^{2})$ Gardim:2016nrr , and ALICE is below ATLAS for most centralities. Interestingly, ATLAS observes a variation of $nMHC(v_{2}^{2},v_{3}^{2})$ going in the same direction as ALICE (down for centralities $40\%$, up for centralities $40\%$) when only particles with $p_{T}>2$ GeV/c/c are included. Since these high-$p_{T}$ particles are likely to have sizable nonflow correlations from jet production, it is tempting to postulate that the difference between ATLAS and ALICE may be due to larger nonflow effects in ALICE, due to the smaller acceptance in pseudorapidity. We consider it likely that ATLAS results are more precise than ALICE results, and we will take this as an excuse for not understanding precisely the ALICE results on higher-order cumulants in Sec. V.
@@
Figure 2: Mixed harmonic cumulants as a function of centrality for Pb+Pb collisions at 5.02 TeV per nucleon pair. The normalized cumulants (5) are taken from Ref. ALICE:2021adw , and the moments in the denominator are evaluated using earlier ALICE data ALICE:2016ccg and formulas in Appendix A. For the negative cumulants, we change the sign before taking the logarithm. $MHC(v_{2}^{4},v_{3}^{4})$ and $MHC(v_{2}^{2},v_{3}^{6})$ are positive for the most central bins, and the three corresponding data points are circled.
@@
For our analysis, we will need the un-normalized cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$. In order to compute them, we evaluate the moments appearing in the denominator of Eq. (5) using standard formulas which are recalled in Appendix A. Results are displayed in Fig. 2 as a function of the collision centrality. The lowest-order cumulant defined by Eq. (4) is negative, as shown in Fig. 1, and the sign alternates as $m$ or $q$ increases, as illustrated by the legend of the figure.
@@ III.1 Definitions and general properties
The central limit theorem implies that only cumulants of order $1$ and $2$ are kept, which means that the probability distribution of $V_{2}$ and $V_{3}$ is Gaussian in the intrinsic frame Voloshin:2007pc . Since none of these cumulants mix $V_{2}$ and $V_{3}$, they are uncorrelated in the Gaussian limit Alqahtani:2025wan . Therefore, the correlations studied in this paper are all driven by non-Gaussian properties of the distribution, so that we shall not neglect cumulants of order $3$ and higher. [^7] Non-Gaussian fluctuations are also responsible for the non-zero $v_{3}\{4\}$ seen in Pb+Pb ALICE:2011ab and Xe+Xe CMS:2019cyz ; CMS:2025xtt collisions , the hierarchy of higher-order cumulants of $v_{2}$ ($v_{2}\{4\}$, $v_{2}\{6\}$, $v_{2}\{8\}$) in p+Pb collisions CMS:2015yux ; CMS:2019wiy , as well as the splitting between $v_{2}\{4\}$ and $v_{2}\{6\}$ in mid-central Pb+Pb collisions Giacalone:2016eyu ; CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg . In Sec. III.2, we discuss the orders of magnitude of these higher-order cumulants and, more specifically, how they depend on the size and shape of the quark-gluon plasma created in the early stages of the collision.
@@ III.2 Power-counting scheme
We now derive general scaling laws that apply to the cumulants in the intrinsic frame. For simplicity, we neglect centrality fluctuations, and we assume that all events in a centrality class have the same impact parameter. The ALICE experiment uses wide centrality bins where this approximation may not be good enough, but this can easily be improved in future analyses. Centrality fluctuations always matter in the 0–5% centrality window Alqahtani:2024ejg , but we have found that their effect on correlations between $v_{2}$ and $v_{3}$ is modest beyond 5% centrality Alqahtani:2025wan , provided that the analysis uses narrow centrality bins ATLAS:2019peb , and we hope that the same conclusion applies to the higher-order cumulants studied in this paper.
@@ IV Relations between experimental cumulants and intrinsic cumulants
We now list the expressions of the higher-order cumulants measured by ALICE ALICE:2021adw , which we truncate by keeping only the leading terms in the power-series expansion:
@@
They are respectively of order ${\cal V}^{8}$, ${\cal V}^{10}$, ${\cal V}^{10}$, ${\cal V}^{12}$, ${\cal V}^{14}$. More generally, $MHC(v_{2}^{2m},v_{3}^{2q})$ is of order ${\cal V}^{2m+4q}$. These orders of magnitude are reflected in the hierarchy observed in Fig. 2. In particular, they explain why $MHC(v_{2}^{6},v_{3}^{2})$ and $MHC(v_{2}^{2},v_{3}^{4})$ are of comparable magnitude, despite being cumulants of different orders (8 and 6 respectively). Since $v_{2}$ and $v_{3}$ are both of order $\cal V$}, the normalized symmetric cumulant (5) is of order ${\cal V}^{2q}$. This explains why the magnitude of $nMHC(v_{2}^{2m},v_{3}^{2q})$ decreases strongly as $q$ increases, as pointed out in Sec. II.
@@
A first comment on the expressions (12), (13) and (14) is that they only involve the mean elliptic flow in the reaction plane, $\bar{V}_{2}$, and mixed cumulants of order $\geq 3$, which quantify non-Gaussian fluctuations. A second remark is that for a given value of $q$, $MHC(v_{2}^{2},v_{3}^{2q})$ contains more terms than higher-order cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ with $m\geq 2$, whose expressions are somewhat simpler. More specifically, the expressions for $m=1$ contain the cumulant $c_{11qq}$ (note that it is the only term which does not vanish for central collisions), which no longer appears for $m\geq 2$. For a given value of $q$, each of the remaining terms is multiplied by a factor proportional to $\bar{V}_{2}^{2}$ as $m$ increases. This can be related to the well-known fact that cumulants of $v_{2}$ of order 4 and higher only depend on $\bar{V}_{2}$ to leading order in $\cal V$}, as recalled in Appendix A. In addition, there are overall numerical factors which increase rapidly as a function of $m$ (22, 44, $24$ for $m=1,2,3$). These numerical factors explain why the normalized cumulants $nMHC(v_{2}^{2m},v_{3}^{2q})$ increase as a function of $m$ for fixed $q$, in absolute magnitude ALICE:2021adw .

### v2

@@ Abstract
The ALICE and CMS Collaborations have analyzed a number of cumulants mixing elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), involving up to $8$ particles, in Pb+Pb collisions at the LHC. We unravel an unexpected simplicity in these complex mathematical quantities for collisions at fixed impact parameter. We show that as one increases the order in $v_{2}$, for a given order in $v_{3}$, the changes in the cumulants are solely determined by the mean elliptic flow in the reaction plane, which originates from the almond-shaped geometry of the overlap area between the colliding nuclei. We derive simple analytic relations between cumulants of different orders on this basis. These relations are in good agreement with recent data from the CMS Collaboration. We argue that agreement will be further improved if the analysis is repeated with a finer centrality binning. We make quantitative predictions for cumulants of order 10 which have not yet been analyzed.
@@ I Introduction
In this paper, we unravel the information contained in cumulants of azimuthal correlations involving both elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), which are called “mixed harmonic cumulants” (MHC) ALICE:2021adw ; CMS:2025opi . $v_{2}$ and $v_{3}$ are the largest Fourier harmonics of anisotropic flow ALICE:2011ab . Relativistic hydrodynamics predicts Niemi:2012aj that they are approximately proportional to the initial anisotropies $\varepsilon_{2}$ and $\varepsilon_{3}$, which are Fourier coefficients of the initial density profile. Therefore, cumulants involving $v_{2}$ and $v_{3}$ constrain models of the initial state more directly than higher harmonics, which are largely driven by nonlinear response terms Borghini:2005kd ; Gardim:2011xv . The lowest-order MHC is a 4-particle cumulant introduced by Bilandzic et al. Bilandzic:2013kga , which has been measured in Pb+Pb collisions at the LHC ALICE:2016kpq ; CMS:2025opi ; ATLAS:2019peb . It represents the linear correlation between $v_{2}^{2}$ and $v_{3}^{2}$, which is negative except in ultra-central collisions. We have recently shown Alqahtani:2025wan that it is driven by the correlation between the elliptic flow in the reaction plane, $v_{2,x}$, and $v_{3}^{2}$. Our goal is to extend this study to higher-order cumulants, of 6 and 8 particles, which have subsequently been measured ALICE:2021adw ; CMS:2025opi .
@@
In Sec. II, we define the cumulants of $v_{2}$ and $v_{3}$ using the formalism of generating functions and we compare data from several analyses in Pb+Pb collisions at the LHC. In Sec. III, we define similar cumulants but in a different coordinate frame, the “intrinsic” frame where the direction of impact parameter is fixed Roubertie:2025qps ; Alqahtani:2025wan . In this frame, the magnitude of cumulants decreases as a function of the order, in a way which follows specific scaling rules, provided that all events have the same impact parameter. In Sec. IV, we express the experimental cumulants as a function of the cumulants in the intrinsic frame. In Sec. V, we derive relations between cumulants of different orders, and compare them with experimental data.
@@ II Cumulants of flow fluctuations in the laboratory frame
II.1 Definitions
The quantities that can be measured experimentally are moments of the joint distribution of $V_{2}$ and $V_{3}$ Bhalerao:2011yg ; Bhalerao:2014xra . The cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ analyzed by ALICE ALICE:2021adw and CMS CMS:2025opi , where $m$ and $q$ are positive integers, are combinations of these moments Bilandzic:2010jr ; Bilandzic:2013kga . [^1] The ALICE collaboration has subsequently analyzed a different set of “cumulants” which are different combination of moments ALICE:2023lwx . They are not derived from the same generating function as the cumulants studied in this paper, and our analysis does not apply to them. Their expressions are derived by expanding the generating function of cumulants Taghavi:2020gcy , which depends on complex variables Mehrabpour:2020wlu $\lambda$ and $\mu$ and their complex conjugates $\lambda^{*}$ and $\mu^{*}$:
@@ II.1 Definitions
It was measured by ALICE in 2016 in Ref. ALICE:2016kpq , where it was named $SC(3,2)$. Higher-order cumulants with $(m,q)=(2,1),(3,1),(1,2),(2,2),(1,3)$ were subsequently measured in Ref. ALICE:2021adw , where their expressions in terms of moments are provided. Deriving these expressions is straightforward using Eq. (2). We do not repeat them because they are lengthy and will not be needed in this work. [^3] Note that the six MHC cumulants analyzed by ALICE and CMS only involve the magnitudes, not the phases of $V_{2}$ and $V_{3}$. The phases only enter at the next order $(m,q)=(3,2)$, a cumulant of order 10 whose expression involves moments of the type $\langle V_{2}^{3}(V_{3}^{*})^{2}\rangle$, and which has not yet been analyzed.
The cumulants are usually normalized as follows:
@@
This normalization suppresses the sensitivity to kinematic cuts: If $v_{n}$ is proportional to the initial anisotropy $\varepsilon_{n}$, the dependence on kinematic cuts can only be in the proportionality factor (the linear response coefficient) which cancels in the ratio (5). The normalized coefficient provides an intuitive dimensionless measure of the strength of the correlation. A value of order unity implies a strong correlation between $v_{2}$ and $v_{3}$.
Figure 1: Centrality dependence of $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb-Pb collisions at 5.02 TeV: Comparison between 2021 ALICE:2021adw and 2024 ALICE:2024nqd ALICE results, ATLAS ATLAS:2019peb and CMS CMS:2025opi . Note that this quantity is denoted by $NSC(3,2)$ in the 2024 ALICE paper, and by $nsc_{2,3}\{4\}$ in the ATLAS paper.
II.2 Experimental data
There have been several analyses of the lowest-order cumulant $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb+Pb collisions at the LHC, which we first compare. We then compare ALICE and CMS results for the higher-order cumulants.
The normalization (5) facilitates the comparison between experiments, because of the reduced sensitivity to kinematic cuts. Fig. 1 displays four sets of experimental results for $nMHC(v_{2}^{2},v_{3}^{2})$ obtained by the ALICE Collaboration in 2021 ALICE:2021adw and 2024 ALICE:2024nqd , by ATLAS ATLAS:2019peb , and recently by CMS CMS:2025opi . One notices that the 2024 ALICE results differ significantly from the 2021 results between 10% and 30% centrality. The 2024 analysis is done in narrower bins (5% as opposed to 10%). The sensitivity of this specific observable to the width of the centrality bins has already been pointed out Gardim:2016nrr , and narrower bins are preferred. Results from ATLAS, CMS, and ALICE 2024 are in fair agreement. The origin of the residual discrepancies is unknown. It could be due to the kinematic cuts, which are different for the four analyses. Note that the dependence on $p_{T}$ cuts cannot be explained within the simple picture of linear hydrodynamic response. The negative sign of $nMHC(v_{2}^{2},v_{3}^{2})$ and the centrality dependence can be ascribed to non-Gaussian fluctuations in the initial state Alqahtani:2025wan , as will be recalled in Sec. IV.
Figure 2: Comparison between ALICE ALICE:2021adw and CMS CMS:2025opi results for higher-order normalized mixed harmonic cumulants.
In this paper, we study higher-order cumulants which were measured in by ALICE in 2021 ALICE:2021adw and more recently by CMS CMS:2025opi . The two sets of results are compared in Fig. 2 as a function of the collision centrality. The normalized cumulants are, in absolute magnitude, smaller than unity, which means that correlations between $v_{2}$ and $v_{3}$ are not strong. [^4] By contrast, $nMHC(v_{2}^{2},v_{4}^{2})$ reaches $1$ in peripheral collisions ATLAS:2019peb ; ALICE:2021adw , which is a natural consequence of the non-linear coupling between $v_{2}$ and $v_{4}$ Teaney:2012ke . The sign of $nMHC(v_{2}^{2m},v_{3}^{2q})$ alternates as $m$ or $q$ increases. In absolute magnitude, it increases somewhat with $m$, and decreases strongly with $q$. These features will be explained in Sec. IV.
As already noticed for the lowest-order cumulant in Fig. 1, there are sizable differences between ALICE and CMS results. CMS uses 5% centrality bins between 10% and 40% centrality, as opposed to 10% bins for ALICE, which may explain some of the discrepancies in this range. The CMS cumulants increase monotonically with the centrality fraction in absolute magnitude in agreement with hydrodynamics predictions shown in ALICE:2021adw ; CMS:2025opi , while the ALICE cumulants reach a maximum and then decrease, and we do not know the origin of this difference.
@@ II.2 Experimental data
Figure 3: Mixed harmonic cumulants as a function of centrality for Pb+Pb collisions at 5.02 TeV per nucleon pair, using CMS data CMS:2025opi for the normalized cumulants and evaluating the denominator of (5) with formulas in Appendix A.
@@
For our analysis, we will need the un-normalized cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$. In order to obtain them from the normalized cumulants, we evaluate the moments appearing in the denominator of Eq. (5) using standard formulas which are recalled in Appendix A. For the sake of illustration, CMS results are displayed in Fig. 3 as a function of the collision centrality. One sees that they span ten orders of magnitude, a point to which we come below in Sec. IV.
@@ III.1 Definitions and general properties
If fluctuations in the intrinsic frame are Gaussian Voloshin:2007pc , all cumulants vanish beyond order $2$. Now, the cumulants of order $1$ and $2$ listed above do not mix $V_{2}$ and $V_{3}$. Therefore, $V_{2}$ and $V_{3}$ are uncorrelated in the Gaussian limit Alqahtani:2025wan , and the correlations studied in this paper are all driven by non-Gaussian properties of the distribution. In order to model these non-Gaussianities, we need to take into account cumulants of order $3$ and higher. [^7] Non-Gaussian fluctuations are also responsible for the non-zero $v_{3}\{4\}$ seen in Pb+Pb ALICE:2011ab and Xe+Xe CMS:2019cyz ; CMS:2025xtt collisions , the hierarchy of higher-order cumulants of $v_{2}$ ($v_{2}\{4\}$, $v_{2}\{6\}$, $v_{2}\{8\}$) in p+Pb collisions CMS:2015yux ; CMS:2019wiy , as well as the splitting between $v_{2}\{4\}$ and $v_{2}\{6\}$ in mid-central Pb+Pb collisions Giacalone:2016eyu ; CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg . In Sec. III.2, we discuss the orders of magnitude of these higher-order cumulants and, more specifically, how they depend on the size and shape of the quark-gluon plasma created in the early stages of the collision.
@@ III.2 Power-counting scheme
We now derive general scaling laws that apply to the cumulants in the intrinsic frame. For simplicity, we neglect centrality fluctuations, and we assume that all events in a centrality class have the same impact parameter. Impact parameter fluctuations have large effects for the most central collisions Alqahtani:2024ejg , but we have found that their effect on correlations between $v_{2}$ and $v_{3}$ is modest beyond 5% centrality Alqahtani:2025wan , provided that the analysis uses narrow centrality bins ATLAS:2019peb . The ALICE experiment uses wide centrality bins and this has a significant effect on the results, as discussed in Sec. II.2, but this can easily be improved in future analyses.
@@ IV Relations between experimental cumulants and intrinsic cumulants
We now list the expressions of the higher-order cumulants measured by ALICE and CMS ALICE:2021adw ; CMS:2025opi , which we truncate by keeping only the leading terms in the power-series expansion:
@@
They are respectively of order ${\cal V}^{8}$, ${\cal V}^{10}$, ${\cal V}^{10}$, ${\cal V}^{12}$, ${\cal V}^{14}$. More generally, $MHC(v_{2}^{2m},v_{3}^{2q})$ is of order ${\cal V}^{2m+4q}$. These orders of magnitude are reflected in the hierarchy observed in Fig. 3. In particular, they explain why $MHC(v_{2}^{6},v_{3}^{2})$ and $MHC(v_{2}^{2},v_{3}^{4})$ are of comparable magnitude, despite being cumulants of different orders (8 and 6 respectively).
Since $v_{2}$ and $v_{3}$ are both of order $\cal V$}, the normalized symmetric cumulant (5) is of order ${\cal V}^{2q}$. This explains why the magnitude of $nMHC(v_{2}^{2m},v_{3}^{2q})$ decreases strongly as $q$ increases, as seen in Fig. 2.
@@
A first comment on the expressions (12), (13) and (14) is that they only involve the mean elliptic flow in the reaction plane, $\bar{V}_{2}$, and mixed cumulants of order $\geq 3$, which quantify non-Gaussian fluctuations. Thus the correlations between $v_{2}$ and $v_{3}$ originate from non-Gaussian fluctuations in the initial state Alqahtani:2025wan .
A second remark is that for a given value of $q$, $MHC(v_{2}^{2},v_{3}^{2q})$ contains more terms than higher-order cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ with $m\geq 2$, whose expressions are somewhat simpler. More specifically, the expressions for $m=1$ contain the cumulant $c_{11qq}$ (note that it is the only term which does not vanish for central collisions), which no longer appears for $m\geq 2$. For a given value of $q$, each of the remaining terms is multiplied by a factor proportional to $\bar{V}_{2}^{2}$ as $m$ increases. This can be related to the well-known fact that cumulants of $v_{2}$ of order 4 and higher only depend on $\bar{V}_{2}$ to leading order in $\cal V$}, as recalled in Appendix A. In addition, there are overall numerical factors which increase rapidly as a function of $m$ (22, 4-4, $24$, 264-264 for $m=1,2,3,4$). These numerical factors explain why the normalized cumulant $nMHC(v_{2}^{2m},v_{3}^{2q})$ changes sign and increases in absolute magnitude as $m$ increases, as seen in Fig. 2.